There are 44 students in a canteen. If there are 28 more girls than boys, how many boys and how many girls are in the canteen?
Answer by nerdybill(7384)

There are 44 students in a canteen. If there are 28 more girls than boys, how many boys and how many girls are in the canteen?
.
Let B = number of boys in the canteen
B+28 = number of girls in the canteen
.
Then, from: "There are 44 students in a canteen." we get our equation:
B + (B+28) = 44
2B + 28 = 44
2B = 44 - 28
2B = 16
B = 8 (number of boys in the canteen)
.
B+28 = 8+28 = 36 (number of girls in the canteen)